OpenStudy (anonymous):

What are the solutions to this inequality ? 8n-14 less than or equal to 13n + 6

OpenStudy (anonymous):

8n-14<= 13n+6 8n-13n - 14 <= 13n - 13n +6 -5n -14+14 <= 6+14 -5n <= 20 -5 n/ -5 >= 20 / -5 (on dividing or multiplying by a -ve number the sense of inequality reverses) n>= - 4 do you understand ?